OpenStudy (anonymous):

Algebra 1 question: AWARDING MEDAL!!

OpenStudy (anonymous):

the linear equation 2x -y = 4 is in standard form . transform the equation to slope intercept form.

OpenStudy (anonymous):

@TheSmartOne @dan815

TheSmartOne (thesmartone):

Do you know what slope-intercept form is? @purpleowl

OpenStudy (anonymous):

mx + b?

TheSmartOne (thesmartone):

\(\sf y=mx+b\)

TheSmartOne (thesmartone):

The y is important :P

OpenStudy (anonymous):

Yes haha

TheSmartOne (thesmartone):

So we have \(\sf 2x -y = 4\) and we need to make it into the form of \(\sf y=mx+b\)

TheSmartOne (thesmartone):

There are 2 ways we can do this. We can subtract 2x from both sides. And then divide -1 from both sides. Or we can add y to both sides. And then subtract 4 from both sides.

OpenStudy (anonymous):

whats -2x minus -1

OpenStudy (anonymous):

-3x?

TheSmartOne (thesmartone):

Nope/

OpenStudy (anonymous):

Idk

OpenStudy (anonymous):

@TheSmartOne there?

TheSmartOne (thesmartone):

Yes, so which method do you want to go with? 1 or 2? 2 is easier I guess :P

OpenStudy (anonymous):

Ok.

OpenStudy (anonymous):

actually 1

TheSmartOne (thesmartone):

So lets see.

OpenStudy (anonymous):

maybe i'll just ask my teacher tomorrow. this is taking too long and i need to get to another things. sorry.

TheSmartOne (thesmartone):

Oh ok then :)